OpenStudy (anonymous):

The volume of a square based pyramid is 81 cubic meters. What is the volume of the cube with the same side length and height? A)9m3 B)27m3 C)243m3 D)729m3 moderator help

OpenStudy (amistre64):

the volume of the base?

OpenStudy (amistre64):

read it wrong

OpenStudy (anonymous):

what

OpenStudy (amistre64):

The volume of "a square based pyramid" is 81 cubic meters. What is the volume of "the cube" with the same side length and height? thats better

OpenStudy (anonymous):

thank you ;)

OpenStudy (amistre64):

the volume of any pointy shape is: (1/3) (base area) (height) to wit; 81 = (1/3) (base area) (h) ; multiply thru by 3 3(81) = s^2 * h ; s^2 refers to the square base area, divide out the h. 243/h = s^2 ; now square root each part sqrt(243/h) = s

OpenStudy (anonymous):

so C)243m3

OpenStudy (amistre64):

hmmm, prolly still not reading it correctly; cube tends to mean a shape with sides and height being equal, but this might mean a cube like shape in general

OpenStudy (amistre64):

if height = side length we get sqrt(243/s) = s 243/s = s^2 243 = s^3 s = cbrt(243) .... might be a better interpretation but its hard to tell with the way it is written

OpenStudy (amistre64):

the volume would then be cbrt(243)^3 = 243m^3 which is prolly the right answer

OpenStudy (anonymous):

so C)243m3